OpenStudy (anonymous):

(6.0x10^7)/(3.0x10^2) What are the steps to solve this and write the answer in scientific notation?

OpenStudy (anonymous):

\[\frac{ (6.0 x 10^7) }{ (3.0 x 10^8) }\]

OpenStudy (anonymous):

How do I do this?

OpenStudy (anonymous):

So, the best way to do this is to get the exponent to be the same. You can do that by moving the decimal on the numerator to the left one place. So you get (.6*10^8)/(3.0*10^8). Then, you can just divide the first parts. So .6/3.0. That gives you 0.2. The last parts cancel out. So the answer is just 0.2.

OpenStudy (anonymous):

If the answer needs to be in scientific notation, it's 2.0*10^1.

OpenStudy (anonymous):

I find it easier this way \[ \frac{6.0\times10^7}{1.0\times10^8}=\frac{6.0}{3.0}\times\frac{10^7}{10^8}= 2\times10^{7-8}=2\times10^{-1} \] key-> keep the exponentials and the numbers separate..

OpenStudy (anonymous):

^-1? I probably messed up somewhere in there.

OpenStudy (anonymous):

The answer key says its 2.0 x 10^5 not 2 x 10^-1

OpenStudy (anonymous):

then you got the question wrong!! \[{6.0\times10^7\over3.0}\times10^8\] this thiswhat the question says?

OpenStudy (anonymous):

The question is (6.0x10^7)/(3.0x10^2) sorry I did write it wrong

OpenStudy (anonymous):

then fowllow the same thing that I did and you'd get the answer

OpenStudy (anonymous):

remember\[{a^m\over a^n}=a^{m-n}\]

OpenStudy (anonymous):

Alright I got it, thanks for the help electrokid